# RealMAN: A Real-Recorded and Annotated Microphone Array Dataset for Dynamic Speech Enhancement and Localization

### v1

@@ Abstract
The training of deep learning-based multichannel speech enhancement and source localization systems relies heavily on the simulation of room impulse response and multichannel diffuse noise, due to the lack of large-scale real-recorded datasets. However, the acoustic mismatch between simulated and real-world data could degrade the model performance when applying in real-world scenarios. To bridge this simulation-to-real gap, this paper presents a new relatively large-scale Real-recorded and annotated Microphone Array speech&Noise (RealMAN) dataset [^1] The RealMAN dataset is publicly available at https://github.com/Audio-WestlakeU/RealMAN.. The proposed dataset is valuable in two aspects: 1) benchmarking speech enhancement and localization algorithms in real scenarios; 2) offering a substantial amount of real-world training data for potentially improving the performance of real-world applications. Specifically, a 32-channel array with high-fidelity microphones is used for recording. A loudspeaker is used for playing source speech signals. A total of 83-hour speech signals (48 hours for static speaker and 35 hours for moving speaker) are recorded in 32 different scenes, and 144 hours of background noise are recorded in 31 different scenes. Both speech and noise recording scenes cover various common indoor, outdoor, semi-outdoor and transportation environments, which enables the training of general-purpose speech enhancement and source localization networks. To obtain the task-specific annotations, the azimuth angle of the loudspeaker is annotated with an omni-direction fisheye camera by automatically detecting the loudspeaker. The direct-path signal is set as the target clean speech for speech enhancement, which is obtained by filtering the source speech signal with an estimated direct-path propagation filter. Baseline experiments demonstrate that i) compared to using simulated data, the proposed dataset is indeed able to train better speech enhancement and source localization networks; ii) using various sub-arrays of the proposed 32-channel microphone array can successfully train variable-array networks that can be directly used to unseen arrays.
@@ 1 Introduction
Microphone array-based multichannel speech enhancement and source localization are two important front-end audio signal processing tasks [1, 2, 3]. Currently, most existing works are deep learning-based and data-driven, for which the amount and diversity of microphone array data are crucial for the proper training of neural networks. Although an infinite amount of microphone array data can be simulated using room impulse response (RIR) simulator (with the image method [4]) and multichannel diffuse noise generator [5], there are still significant mismatches between the acoustic properties of the simulated and real-world microphone array data, including 1) The geometry of the room and the presence of obstacles within it. The image method [4] simulates empty box-shaped rooms, which does not accurately reflect the acoustics of the real-world rooms with irregular shapes and built-in obstacles. 2) The directivity of the sound sources and microphones, as well as the wall absorption coefficients [6]. 3) Piece-wise simulation of moving sound source [7]. To simulate microphone signals from the moving source, the continuous trajectory of the sound source is discretized into multiple locations. One signal segment is simulated for each location, and signal segments are then connected. If the sound source moves too fast or the discretization is too sparse, the resulting microphone signals may contain clicking noises and other audible artifacts [7]. 4) The spatial correlation of the simulated multi-channel noise is normally determined under the hypothesis of a theoretical diffuse noise field [5], while the real-world noises could deviate from the theoretical values by a large margin, as shown in Fig. 5. Relevant studies have shown that models trained on simulated data often perform poorly in real-world scenarios [6, 8, 9].
@@
Training with real-world data can avoid these mismatches. However, existing real-world microphone array data with annotated target clean speech and source location information is limited and lacks diversity. To address this, we collect a new Real-recorded and annotated Microphone Array speech&Noise (RealMAN) dataset from a variety of real-world indoor, outdoor, semi-outdoor, and transportation scenes. These recordings encompass a diverse range of spatial/room acoustics and noise characteristics. The dataset consists of 83 hours of speech and 144 hours of noise, recorded in 32 and 31 different scenes respectively, with both speech and noise recorded in 17 of these scenes. Both static and moving speech sources are included. We provide annotations of source azimuth angle, direct-path target clean speech, and speech transcription, for speech enhancement (and evaluation of automatic speech recognition) and source localization. A 32-channel microphone array is used for recording. End-to-end speech enhancement and source localization models are normally array-dependent, which means that the network trained with one specific array can be only used for the same array. However, collecting real-world data for a new array is cumbersome. One solution is to use data from various arrays to train a variable-array network that can generalize to unseen arrays [10, 11, 12]. Our 32-channel array can provide many different sub-arrays for training such variable-array networks. Baseline experiments have been conducted on the proposed dataset, demonstrating that 1) Comparing with the simulated data, training with the real-world data eliminates the simulation-to-real problem and achieves better performances in speech enhancement and source localization. Thus, the proposed dataset is more suitable for benchmarking new algorithms and reflecting their capabilities; 2) the variable-array networks [10, 12] can be successfully trained with our 32-channel array dataset. Hopefully, these networks can be applied directly to real applications involving unseen arrays.
@@ 2 Related Work
Multi-channel speech recording and annotation. In MIR [13], BUTReverb [14], Reverb [15], DCASE [16], ACE [17] and dEchorate [18], real-world RIRs are measured instead of directly collecting speech recordings. Measuring real-world RIRs offers several advantages: 1) speech recordings can be generated by convolving RIRs with source signals, which provides sufficient data realness; 2) The direct-path signal can be easily obtained by convolving the direct-path impulse response (extracted from RIRs) with source signals, which can be used as the training target signal for speech enhancement; 3) information for source localization, such as the time-difference of arrival, can also be obtained from the multi-channel direct-path impulse responses. However, one significant drawback of RIR measurement is that it is more time-consuming than speech recording. Consequently, existing datasets such as MIR, BUTReverb, Reverb, ACE, and dEchorate offer only a limited range of measured RIRs and scenes. The scenario becomes even more complex with a moving source, as seen in DCASE, which requires measuring RIRs at multiple discrete locations along a single trajectory, making the process substantially more time-consuming.
Other datasets directly provide multi-channel speech recordings. However, annotating the direct-path speech (as the training target signal) for speech enhancement poses challenges. Although some datasets provide close-talking signals, these cannot serve as the target signals since the direct-path speech is essentially an energy-attenuated and time-shifted version of the close-talking signals. To obtain a clean target signal, the CHiME-3 dataset [19] simulates the time delay of the direct-path speech for training, and also provides the speech signals recorded in a booth for development and test. In addition to evaluating the speech quality, speech enhancement can also be evaluated in terms of automatic speech recognition (ASR) performance. LibriCSS [20], MC-WSJ-AV [21], CHiME-5/-6/-7 [22], AMIMeeting [23], AISHELL-4 [24] and AliMeeting [25] (see Appendix for details) provide speech transcriptions for evaluating the ASR performance. Due to the lack of target signal, the speech enhancement network for these datasets can only be trained with simulated data.
DCASE [16] and LOCATA [26] provide the source location. In DCASE, the RIR needs to be finely measured at many pre-defined discrete locations, which is very time-consuming. In LOCATA, source location is obtained through an optical tracking system installed in the computing laboratory. However, setting up the optical tracking system for multiple rooms incurs significant costs.
@@
# Scenes Scene type Source state # RIR / speech duration
MIR [13] 3 Lab Static 78 RIR 8-ch linear ($\times$3)
BUTReverb [14] 9 - Static 51 RIR 8-ch spherical
Reverb [15] 3 - Static 24 RIR 8-ch circular
DCASE [16] 9 Campus Static 38530 RIR, Location 32-ch spherical
ACE [17] 7 Campus Static 14 RIR, RoomInfo 2-ch, 3-ch triangle, 8-ch linear, 5-ch cruciform, 32-ch spherical
dEchorate [18] 11 Lab Static 99 RIR, RoomInfo 5-ch linear ($\times$6)
CHiME-3/-4 [19] 5 Multiple - 9.9 h Recording, Transcription 6-ch rectangular
LOCATA [26] 1 Lab Static+moving 0.9 h Recording, Location 15-ch planar, 32-ch spherical, 12-ch robot, 4-ch hearing aids
RealMAN (prop.) 32 Multiple Static+moving 83 h Recording, Transcription, Direct-path signal, Location 32-ch (include various sub-arrays)
@@
Dataset Noise Scenes / Types Duration
BUTReverb [14] 9 rooms (large, middle and small size), with room environmental noise (silence) 4.7 h
Reverb [15] 3 rooms (large, medium and small), with stationary background noise mainly caused by air conditioning 0.8 h
ACE [17] 7 offices, with meeting and teaching rooms, babble, ambient and fan noise recorded in each room 13.6 h
dEchorate [18] 1 room with 11 surface absorptions, with diffuse babble, white and silence noises 0.6 h
CHiME-3/-4 [19] 4 public scenarios (cafe, street junction, public transport and pedestrian area) 8.4 h
DCASE [16] 9 rooms in university buildings, with ambient noise 3.9 h
DEMAND [27] 18 scenes (domestic, office, public, transportation, nature categories) 1.5 h
RealMAN (prop.) 31 scenes (indoor, semi-outdoor, outdoor, transportation categories) 144 h
@@
By comparison, the proposed RealMAN dataset has the following advantages. 1) Realness. Speech and noise are recorded in real environments. Direct recording for moving sources avoids issues associated with the piece-wise generation method. Different individuals move the loudspeaker freely to closely mimic the human walking in real applications. One unreal factor is that we use a loudspeaker playing back speech signals, instead of speaking by real human speakers. 2) Quantity and diversity. We record both speech signals and noise signals across various scenes. Compared with existing datasets, our collection offers greater diversity in spatial acoustics (in terms of acoustic scenes, source positions and states, etc) and noise types. This enables effective training of speech enhancement and source localization networks. 3) Annotation. We provide detailed annotations for source locations, direct-path speech, and speech transcriptions, which are essential for accurate training and evaluation. 4) Number of channels. The number of microphone channels, i.e. 32, is higher than almost all existing datasets, which facilitates the training of variable-array networks. 5) Recording cost. The recording, playback, and camera devices are portable and easily transportable to different scenes. The camera shares a common coordinate center with the microphone array, simplifying the annotation process of source locations.
@@ 3.1 Recording system
32-channel microphone array is comprised of 32 high-fidelity Audio-Technica BP899 microphones. The array geometry is shown in Fig. 1(b). This array encompasses the array topology found in common use cases, including common planar linear arrays, circular arrays, and 3D arrays. The sampling rate of microphone recording is 48 kHz. The sampled audio signals are then digitized by 4 clock-synchronized 8-channel microphone pre-amplifiers (RME OctoMic II) and processed by a laptop through an audio interface (Digiface USB).
@@
High-fidelity monophonic speaker: A high-fidelity monophonic loudspeaker (FOSTEX 6301 NE) is used to play source speech signals. It is placed on a height-adjustable and mobile carrier such that one can control the position of the loudspeaker to mimic a standing/moving human speaker. A 5-cm diameter LED light is put on the top of the loudspeaker to magnify the visibility of loudspeaker to the the fisheye camera and annotate the position of the loudspeaker. The LED light can emit red or green light, which is visible for the fisheye camera under various of light conditions.
@@ 3.2 Source speech signals
Source speech signals that are played by the loudspeaker contains nearly 35 hours of clean Mandarin speech, of which about 30 hours are free-talking and 5 hours are reading. For free talk, speakers are encouraged to converse alone. Reading speech entail speakers reading news articles. The topics of speech content spread a wide range of domains including news reports, games, reading experiences, and life trivia. There are 55 speakers in total, including 27 males and 28 females. 17 speakers are recorded in a studio (T60 is smaller than 150 ms) with a high-fidelity microphone, while the rest are recorded in a living room with a bit larger reverberation time (T60 is about 200 ms) using a lower-fidelity microphone. The distance between speaker and microphone is about 0.2 m.
@@ 3.3 Speech and noise recording process
Speech. The objective of the speech recording process is to mirror real-life scenarios of human activities. In each scene, the position of both the camera and microphone array are fixed. When playing source speech, the position of the loudspeaker takes on either static or moving states. For the moving case, one person manually moves the loudspeaker carrier with varying but reasonable moving speed. In transportation scenarios, people typically maintain a stationary position, thereby the loudspeaker only takes the static state. The height of the microphone array is set to 1.40 m. The center height of the loudspeaker is aligned with the height of the mouth of a standing person, varying randomly between 1.30 m and 1.60 m. Most of the time, the loudspeaker faces towards the microphone array. Speech data was recorded across 32 distinct environments, including indoor, outdoor, semi-outdoor, and transportation scenarios. The detailed scene information of the proposed dataset is shown in Appendix B.1. We ensure that most speech recordings were conducted under quiet conditions (usually at midnight), with background noise levels maintained below 40 dB.
Noise. Noise recording is simpler, for which we place the microphone array in various environments to capture the real-world ambient noise. Noise recording is normally conducted in the daytime with active events in each environment. The collected recording clips with noise power lower than a certain threshold are abandoned. Then, an advanced voice activity detection is conducted to further filter out those recording clips including prominent speech signals. Noise is recorded in 31 different scenarios and ultimately retained 144.45 hours of recordings, covering most everyday scenarios.

### v2

@@ Abstract
The training of deep learning-based multichannel speech enhancement and source localization systems relies heavily on the simulation of room impulse response and multichannel diffuse noise, due to the lack of large-scale real-recorded datasets. However, the acoustic mismatch between simulated and real-world data could degrade the model performance when applying in real-world scenarios. To bridge this simulation-to-real gap, this paper presents a new relatively large-scale Real-recorded and annotated Microphone Array speech&Noise (RealMAN) dataset [^1] The RealMAN dataset is publicly available at https://github.com/Audio-WestlakeU/RealMAN.. The proposed dataset is valuable in two aspects: 1) benchmarking speech enhancement and localization algorithms in real scenarios; 2) offering a substantial amount of real-world training data for potentially improving the performance of real-world applications. Specifically, a 32-channel array with high-fidelity microphones is used for recording. A loudspeaker is used for playing source speech signals (about 35 hours of Mandarin speech). A total of 83.7 hours of speech signals (about 48.3 hours for static speaker and 35.4 hours for moving speaker) are recorded in 32 different scenes, and 144.5 hours of background noise are recorded in 31 different scenes. Both speech and noise recording scenes cover various common indoor, outdoor, semi-outdoor and transportation environments, which enables the training of general-purpose speech enhancement and source localization networks. To obtain the task-specific annotations, speaker location is annotated with an omni-directional fisheye camera by automatically detecting the loudspeaker. The direct-path signal is set as the target clean speech for speech enhancement, which is obtained by filtering the source speech signal with an estimated direct-path propagation filter. Baseline experiments demonstrate that i) compared to using simulated data, the proposed dataset is indeed able to train better speech enhancement and source localization networks; ii) using various sub-arrays of the proposed 32-channel microphone array can successfully train variable-array networks that can be directly used to unseen arrays.
@@ 1 Introduction
Microphone array-based multichannel speech enhancement and source localization are two important front-end audio signal processing tasks [1, 2, 3]. Currently, most existing works are deep learning-based and data-driven, for which the amount and diversity of microphone array data are crucial for the proper training of neural networks. Although an unlimited amount of microphone array data can be generated using simulated room impulse response (RIR) (with the image source method, ISM [4]) and multichannel diffuse noise [5], there are still significant mismatches between the acoustic properties of the simulated and real-world data, including 1) The directivity of microphones and sound sources, as well as the wall absorption coefficients. ISM [4] assumes both microphones and sound sources to be omni-directional, while real microphones (especially when mounted on specific devices) have a certain directivity and human speakers have a frontal (on-axis) directivity [6]. In addition, ISM normally uses frequency-independent wall absorption coefficients, which violates the case of real walls. In [7], the authors introduced the frequency-dependent microphone directivity, sound source directivity and wall absorption coefficients into ISM, which largely improves the realness of simulated RIRs. 2) Room geometry and furniture. ISM normally simulates empty shoebox-shaped rooms, while real rooms could have irregular geometry and built-in furniture. For simulating irregular rooms and built-in furniture, more advanced (but less computationally efficient) geometric acoustic (ray-tracing) methods can be used [8]. One extra difficulty is to configure the room geometry and layout, room and furniture materials to conform to the real-world scene semantics. In [8], this problem is resolved by leveraging scene computer aided design (CAD) models for room design and natural language processing techniques for material setup, which largely improves the realness of simulated RIRs as well. 3) Piece-wise simulation of moving sound source [9]. To simulate microphone signals from a moving sound source, the continuous trajectory of the source must be discretized into densely-distributed locations. One signal segment is simulated for each location, and signal segments are then connected. If the sound source moves too fast or the discretization is too sparse, the resulting microphone signals may contain clicking noises and other audible artifacts [9]. 4) The spatial correlation of simulated multi-channel noise is normally determined under the hypothesis of a theoretical diffuse noise field [5], while the real-world ambient noise is a superposition of a large number of time-varying (partially) diffuse and directional noise sources and its spatial correlation could largely deviate from the theoretical values. Overall, due to the simulation-to-real mismatch, relevant studies on various tasks have shown that models trained with simulated data often perform poorly in real-world scenarios [7, 10, 11, 12].
@@
Training with real-world data can avoid these mismatches. However, existing real-world microphone array data with annotated target clean speech and source location information is limited and lacks diversity. To address this, we collect a new Real-recorded and annotated Microphone Array speech&Noise (RealMAN) dataset from a variety of real-world indoor, outdoor, semi-outdoor, and transportation scenes. These recordings encompass diverse spatial/room acoustics and noise characteristics. A loudspeaker is used for playing source speech signal (although the used loudspeaker has a similar frontal on-axis directivity as human speakers [6], note that one fixed loudspeaker cannot account for the various directivities across different human speakers), and about 35 hours of clean Mandarin speech are used as source speech (more speech materials and for other languages will be considered in future recording if copyright can be issued). The dataset consists of 83.7 hours of speech and 144.5 hours of noise, recorded in 32 and 31 different scenes respectively, with both speech and noise recorded in 17 of these scenes. Both static and moving speech sources are included. We provide annotations of direct-path target clean speech, speech transcription, and source location (azimuth angle, elevation angle and distance relative to the microphone array), for speech enhancement (and evaluation of automatic speech recognition, ASR) and source localization. Where the direct-path target clean speech is obtained by filtering the source speech with estimated direct-path impulse responses, and the source locations are annotated with an omni-directional fisheye camera. A 32-channel microphone array is used for recording. End-to-end speech enhancement and source localization models are normally array-dependent, namely the network trained with one specific array can be only used for the same array. However, collecting real-world data for every new array is cumbersome. One solution is to use data from various arrays to train a variable-array network that can generalize to unseen arrays [13, 14, 15]. Our 32-channel array can provide many different sub-arrays for training such variable-array networks. Baseline experiments have been conducted on the proposed dataset, demonstrating that 1) Compared with using simulated data, training with the proposed real data eliminates the simulation-to-real problem and achieves better performances in speech enhancement and source localization. Thus, the proposed dataset is more suitable for benchmarking new algorithms and evaluating their actual capabilities; 2) the variable-array networks [13, 15] can be successfully trained with our 32-channel array dataset. Hopefully, these networks can be applied directly to real applications involving unseen arrays.
@@ 2 Related Work
Multi-channel speech recording and annotation. In MIR [16], BUTReverb [17], Reverb [18], DCASE [19], ACE [20] and dEchorate [21], real-world RIRs are measured instead of directly collecting speech recordings. Measuring real-world RIRs offers several advantages: 1) microphone array speech can be generated by convolving RIRs with source signals, which exhibits sufficient data realness; 2) Direct-path speech can be easily obtained by convolving the direct-path impulse response (extracted from RIRs) with source signals, which can be used as the training target signal for speech enhancement; 3) information for source localization, such as the time-difference of arrival, can also be obtained from the multi-channel direct-path impulse responses. However, one significant drawback of RIR measurement is that it is more time-consuming than speech recording. Consequently, existing datasets such as MIR, BUTReverb, Reverb, ACE, and dEchorate offer only a limited range of measured RIRs and scenes. As for moving source, to obtain the RIR at densely-distributed discrete locations along a moving trajectory, the measurement process becomes even more time-consuming.
Other datasets directly provide multi-channel speech recordings. However, annotating the direct-path speech (as the training target signal) for speech enhancement poses challenges. Although some datasets provide close-talking signals, these cannot serve as the target signals since the direct-path speech is essentially an energy-attenuated and time-shifted version of the close-talking signals. To obtain a clean target signal, the CHiME-3 dataset [22] simulates the time delay of the direct-path speech for training, and also provides the speech signals recorded in a booth for development and test. In addition to evaluating the speech quality, speech enhancement can also be evaluated in terms of ASR performance. LibriCSS [23], MC-WSJ-AV [24], CHiME-5/-6/-7 [25], AMIMeeting [26], AISHELL-4 [27] and AliMeeting [28] (see Appendix for details) provide speech transcriptions for evaluating the speech enhancement performance via ASR. Due to the lack of target signal, the speech enhancement network for these datasets can only be trained with simulated data.
The annotation of sound source location is also not easy. VoiceHome-2 [29], LOCATA [30] and STARSS22/23 [31, 32] record speech of real human speakers and provide the speaker location. In VoiceHome-2, speakers can only locate at a small number of pre-defined positions, and their coordinates were measured with a laser telemeter. In LOCATA and STARSS22/23, speaker locations are obtained through an optical tracking system. However, the difficulty of setting up the optical tracking system limits the number of recording scenes.
@@
# scene scene type source state # RIR / speech duration
MIR [16] 3 Lab Static 78 RIRs RIR 8-ch linear ($\times$3)
BUTReverb [17] 9 - Static 51 RIRs RIR 8-ch spherical
Reverb [18] 3 - Static 24 RIRs RIR 8-ch circular
DCASE [19] 9 Campus Static 38530 RIRs RIR, location 32-ch spherical
ACE [20] 7 Campus Static 14 RIRs RIR, roomInfo 2-ch, 3-ch triangle, 8-ch linear, 5-ch cruciform, 32-ch spherical
dEchorate [21] 11 Lab Static 99 RIRs RIR, roomInfo 5-ch linear ($\times$6)
CHiME-3/-4 [22] 5 Multiple - 9.9 hours Recording, transcription 6-ch rectangular
VoiceHome-2 [29] 12 Home Static 5 hours Recording, location, transcription 8-ch
LOCATA [30] 1 Lab Static, moving 0.9 hours Recording, location 15-ch planar, 32-ch spherical, 12-ch robot, 4-ch hearing aids
STARSS22 [31] 16 Campus, corporation Moving 6.9 hours Recording, location, sound event labels 32-ch spherical
STARSS23 [32] 16+ Campus, corporation Moving 10.9 hours Recording, location, sound event labels 32-ch spherical
RealMAN (prop.) 32 Multiple Static+ moving 83.7 hours Recording, location, direct-path signal, transcription 32-ch (include various sub-arrays)
@@
Dataset Noise Scene / Type Duration
BUTReverb [17] 9 rooms (large, middle and small size), with room environmental noise (silence) 4.7 hours
Reverb [18] 3 rooms (large, medium, and small), with stationary background noise mainly caused by air conditioning 0.8 hours
ACE [20] 7 offices, with meeting and teaching rooms, babble, ambient and fan noise recorded in each room 13.6 hours
dEchorate [21] 1 room with 11 surface absorptions, with diffuse babble, white and silence noises 0.6 hours
CHiME-3/-4 [22] 4 public scenarios (cafe, street junction, public transport and pedestrian area) 8.4 hours
DCASE [19] 9 rooms in university buildings, with ambient noise 3.9 hours
DEMAND [33] 18 scenes (domestic, office, public, transportation, nature categories) 1.5 hours
RealMAN (prop.) 31 scenes (indoor, semi-outdoor, outdoor, transportation categories) 144.5 h
@@
By comparison, the proposed RealMAN dataset has the following advantages. 1) Realness. Speech and noise are recorded in real environments. Direct recording for moving sources avoids issues associated with the piece-wise generation method. Different individuals move the loudspeaker freely to closely mimic human movements in real applications. 2) Quantity and diversity. We record both speech signals and noise signals across various scenes. Compared with existing datasets, our collection offers greater diversity in spatial acoustics (in terms of acoustic scenes, source positions and states, etc) and noise types. This enables effective training of speech enhancement and source localization networks. 3) Annotation. We provide detailed annotations for direct-path speech, speech transcriptions and source location, which are essential for accurate training and evaluation. 4) Number of channels. The number of microphone channels, i.e. 32, is higher than almost all existing datasets, which facilitates the training of variable-array networks. 5) Relatively low recording cost. The recording, playback, and camera devices are portable and easily transportable to different scenes.
@@ 3.1 Recording system
32-channel microphone array is comprised of 32 high-fidelity omni-directional Audio-Technica BP899 microphones. The microphone’s frequency response range is 20 Hz to 20 kHz. The array geometry is shown in Fig. 1(b). This array encompasses the array topology found in common use cases, including common planar linear arrays, circular arrays, and 3D arrays. The sampling rate of microphone recording is 48 kHz. The sampled audio signals are then digitized by 4 clock-synchronized 8-channel microphone pre-amplifiers (RME OctoMic II) and processed by a laptop through an audio interface (RME Digiface USB).
@@
High-fidelity monophonic speaker: A high-fidelity monophonic loudspeaker (FOSTEX 6301 NE) is used to play source speech signals. The loudspeaker’s frequency response range is 70 Hz to 15 kHz. It is placed on a height-adjustable and mobile carrier such that one can control the position of the loudspeaker to mimic a standing/moving human speaker. A 5-cm diameter LED light is put on the top of the loudspeaker to magnify the visibility of loudspeaker to the the fisheye camera and annotate the position of the loudspeaker. The LED light can emit red or green light, which is visible for the fisheye camera under various light conditions.
@@ 3.2 Source speech signals
Source speech signals that are played by the loudspeaker contains nearly 35 hours of clean Mandarin speech (close-sourced), of which about 30 hours are free-talking and 5 hours are reading. For free talk, speakers are encouraged to converse alone. Reading speech entail speakers reading news articles. The topics of speech content spread a wide range of domains including news reports, games, reading experiences, and life trivia. There are 55 speakers in total, including 27 males and 28 females. 17 speakers are recorded in a studio (T60 is smaller than 150 ms) with a high-fidelity microphone, while the rest are recorded in a living room with a bit larger reverberation time (T60 is about 200 ms) using a lower-fidelity microphone. The distance between speaker and microphone is about 0.2 m.
@@ 3.3 Speech and noise recording process
Speech. The objective of the speech recording process is to mirror the human activities in real speech applications, such as (multi-party) human-robot interaction, smart loudspeaker, meeting, etc. In each scene, the position of both the camera and microphone array are fixed. When playing source speech, the position of the loudspeaker takes on either static or moving states. For the moving case, one person manually moves the loudspeaker carrier with varying but reasonable moving speed. In transportation scenarios, people typically maintain a stationary position, thereby the loudspeaker only takes the static state. The height of the microphone array is always set to 1.40 m. In each scene, the microphone array is located at one position around the scene center part, and sometimes it is put close (not very close) to one side wall for having a large source-to-array distance or not disturbing other people. The center height of the loudspeaker is aligned with the height of the mouth of a standing person, varying randomly between 1.30 m and 1.60 m. Speech data was recorded across 32 distinct environments, including indoor, outdoor, semi-outdoor, and transportation scenarios. We ensure that most speech recordings were conducted under quiet conditions (usually at midnight). Over all scenes, the sound pressure level (SPL) of speech recordings and silent backgrounds are averagely about 68 dB (61 dBA) and 57 dB (36 dBA), respectively.
In different speech applications, speakers could have very different states in terms of facing or not facing the device, static or moving, small-range pacing or large-range walking, and head turning. For example, as for human-robot interaction and smart loudspeaker, the speaker normally faces the device and could have some movements. In a meeting application, the speakers normally face each other and have head turning. During speech recording, we do consider these factors to a certain extent. For static speaker, most of the time, the loudspeaker faces towards the microphone array, with a small portion of side-facing (but no back-facing). For moving speaker, movements include a large portion of large-range walking, a small portion of small-range pacing and a smaller portion of head turning. The moving trajectory of large-range walking can be axial, tangential or in between, and can be random walking as well. During movement, the loudspeaker orientation is put either facing (most of the time side-facing) the microphone array, or towards the moving direction. The source-to-array distances are mainly distributed in the range of 0.5 m $\sim$ 5 m. Overall, in the proposed dataset, a number of speaker states are considered. However, we think it is still far from exhausting the speaker states in various speech applications, and the influence of speaker states on speech enhancement and source localization would be an interesting topic for future research.
The detailed speech recording information for each scene is given in Appendix B.1, including the recording duration, SPLs, room size and reverberation time. The statistics of source locations and several examples of speaker moving trajectory are given in Appendix B.4.
Noise. Noise recording is simpler, for which we place the microphone array in various environments to capture the real-world ambient noise. Noise recording is normally conducted in the daytime with active events in each environment. The collected recording clips with noise power lower than a certain threshold are abandoned. Then, an advanced voice activity detection is conducted to further filter out those recording clips including prominent speech signals. Noise is recorded in 31 different scenarios and ultimately retained 144.5 hours of recordings, covering most everyday scenarios. The SPL of noise recordings is averagely about 71 dB (58 dBA) over all scenes. The duration and SPLs of noise recording for each scene are given in Appendix B.1 as well.